Use the mean value of 540 for Keq to calculate the SCN- concentration in a solution whose initial Fe3+ concentration was 3.60x10-2 M and initial SCN- concentration was 1.30x10-3 M. Give answer in M.

Equation: Fe3+ + SCN- <-----------> FeNCS2+
Answer #1

Fe^3+ (aq) + SCN^- (aq) ------------> FeNCS^2+(aq)

I           0.03 0.001     0

C          -x                    -x    +x

E         0.03-x            0.001-x +x

         Kc   = [FeNCS^2+]/[Fe^3+][SCN^-]

540 = x/(0.03-x)(0.001-x)

540*(0.03-x)(0.001-x)   = x

       x = 0.000940

   [Fe^3+]    = 0.03-x   = 0.03-0.000940 = 0.02906M

[SCN-]     = 0.001-x   = 0.001-0.000940 = 0.6*10^-4M

   [FeNCS^2+]   = x   = 0.000940M